Consider a hypothetical economy characterized by the following behavioral equation IS                                  LM:

C=400-200r+0.2Y              Md=Y-1000r

I= 240-400r                        Ms=800

G=200                                 P=2

a)     derive the equations of the IS and LM curves of the economy

b)     Find the equilibrium income level and the interest rate simultaneously.



1
Expert's answer
2021-08-16T10:02:11-0400

(a)

C"=400-200r+0.2Y"

I"=240-400r"

G"=200"

Y"=C+I+G"

Y"=" 400"-" 200r"+" 0.2Y"+" 240"-" 400r"+" 200

Y"-0.2Y=" 840"-" 600r

0.8Y"=" 840"-" 600r

LS:

Y"=" 1050"-"750r

LM;

Md"=" Y"-" 1000r

Ms"=" 800

P"=" 2

Ms"=" Mt"+" Mz

800"=" Y"-"1000r"+" 2

800"+" 1000r"-" 2"=" Y

Y"=" 798"+" 1000r

(b)

1050"-" 750r"=" 798"+" 1000r

1050"-" 798"=" 1000r"+" 750r

r"=" 0.144

"=" 14.4%

Substitute r"="0.144

Y"=" 1050"-"750"(0.144)"

Y"=" 1050"-" 108

Y"=" 942